15 Years 65 Years 165 Years 12 15 Douglas-fir 44 WPlease show work, thank you
Answer #1

Simpson's reciprocal of 65 year old stand = 1/D

where D = summation of [ n(n-1)] / N(N-1)

We have to calculate n(n-1) for each species

the values of n(n-1) for different species from Douglas Fir to Wild Cherry are

15*14 , 47*46 , 9*8 , 0 , 0 , 0 , 0 , 3 * 2 , 12*11 , 0

= 210 , 2162 , 72 , 0 , 0 , 0 , 0 , 6 , 132 , 0

summation of n(n-1) = 210+2162+72+6+132 = 2582

Now,

N = summation of n = 15+47+9+1+3+12+1 = 88

So , N(N-1) = 88*87 = 7656

Hence, D = 2582 / 7656 = 0.34

So,

Simpson's Reciprocal index = 1/D = 1/0.34 = 2.9